# FLUX that Plays Music

### v1

@@ 1 Introduction
Music, as a form of artistic expression, holds profound cultural importance and resonates deeply with human experiences (Briot et al., 2017). The task of text-to-music generation, which involves converting textual descriptions of emotions, styles, instruments, and other musical elements into audio, offers innovative tools and new avenues for multimedia creation (Huang et al., 2023b). Recent advancements in generative models have led to significant progress in this area (Yang et al., 2017; Dong et al., 2018; Mittal et al., 2021). Traditionally, approaches to text-to-music generation have relied on either language models or diffusion models to represent quantized waveforms or spectral features (Agostinelli et al., 2023; Lam et al., 2024; Liu et al., 2024; Evans et al., 2024; Schneider et al., 2024; Fei et al., 2024a, 2023c; Chen et al., 2024b). Among these, diffusion models (Song et al., 2020), trained to reverse the process of data transformation from structured states to random noise (Sohl-Dickstein et al., 2015; Song & Ermon, 2020), have shown exceptional effectiveness in modeling high-dimensional perceptual data, including music (Ho et al., 2020; Huang et al., 2023c; Ho et al., 2022; Kong et al., 2020b; Rombach et al., 2022).
@@
Given the iterative nature of diffusion process, coupled with the significant computational costs and extended sampling times during inference, there has been a growing body of research focused on developing more efficient training strategy and accelerating sampling schedule  (Karras et al., 2023; Liu et al., 2022a; Lu et al., 2022a; Fei, 2019; Lu et al., 2022b; Kingma & Gao, 2024), such as distillation (Sauer et al., 2024; Song et al., 2023; Song & Dhariwal, 2023). A particularly effective approach involves defining a forward path from data to noise, which facilitates more efficient training (Ma et al., 2024a) as well as better generative performance. One effective method among them is the Rectified Flow (Liu et al., 2022a; Albergo & Vanden-Eijnden, 2022; Lipman et al., 2023), where data and noise are connected along a linear trajectory. It offers improved theoretical properties and has shown promising results in image generation(Ma et al., 2024a; Esser et al., 2024), however, its application in music creation remains largely unexplored.
@@
In the design of model architectures, traditional diffusion models frequently employ U-Net (Ronneberger et al., 2015) as the foundational structure. However, the inherent inductive biases of convolutional neural networks inadequately captures the spatial correlations within signals (Esser et al., 2021) and are insensitive to scaling laws (Li et al., 2024b). Transformer-based diffusion models have effectively addressed these limitations (Peebles & Xie, 2023; Bao et al., 2023; Fei et al., 2024b, c) by treating images as sequences of concatenated patches and utilizing stacked transformer blocks for noise prediction. The incorporation of cross-attention for integrating textual information has established this approach as the standard for generating high-resolution images and videos from natural language descriptions, demonstrating impressive generalization capabilities (Chen et al., 2023, 2024a; Fei et al., 2024e; Esser et al., 2024; Fei et al., 2023b; Ma et al., 2024b, b; Yang et al., 2024). Notably, the recently open-sourced FLUX model, with its well-designed structure, exhibits strong semantic understanding and produces high-quality images, positioning it as a promising framework for conditional generation tasks.
@@ 2.1 Text-to-music Generation
Text-to-music generation seeks to produce music clips that correspond to descriptive or summarized text inputs. Prior approaches have primarily employed language models (LMs) or diffusion models (DMs) to generate quantized waveform representations or spectral features. For generating discrete representation of waveform, models such as MusicLM (Agostinelli et al., 2023), MusicGen (Copet et al., 2024), MeLoDy (Lam et al., 2024), and JEN-1 (Li et al., 2024c) utilize LMs and DMs on residual codebooks derived from quantization-based audio codecs (Zeghidour et al., 2021; Défossez et al., 2022). Conversely, models like Moûsai (Schneider et al., 2024), Noise2Music (Huang et al., 2023b), Riffusion (Forsgren & Martiros, 2022), AudioLDM 2 (Liu et al., 2024), MusicLDM (Chen et al., 2024b), and StableAudio (Evans et al., 2024) employ U-Net-based diffusion techniques to model mel-spectrograms or latent representations obtained through pretrained VAEs, subsequently converting them into audio waveforms using pretrained vocoders (Kong et al., 2020a). Additionally, models such as Mustango (Melechovsky et al., 2023) and Music Controlnet (Wu et al., 2024) incorporate control signals or personalization (Plitsis et al., 2024; Fei et al., 2023a), including chords and beats, in a manner similar to ControlNet (Zhang et al., 2023). Our method along with this approach by modeling the mel-spectrogram within a latent VAE space.
@@ 2.2 Diffusion Transformers
Transformer architecture (Vaswani et al., 2017) has achieved remarkable success in language models (Radford, 2018; Radford et al., 2019; Raffel et al., 2020) and has also demonstrated significant potential across various computer vision tasks, including image classification (DOSOVITSKIY, 2020; He et al., 2022; Touvron et al., 2021; Zhou et al., 2021; Yuan et al., 2021; Han et al., 2021), object detection (Liu et al., 2021; Wang et al., 2021, 2022; Carion et al., 2020), and semantic segmentation (Zheng et al., 2021; Xie et al., 2021; Strudel et al., 2021), among others (Sun et al., 2020; Li et al., 2022b; Zhao et al., 2021; Liu et al., 2022b; He et al., 2022; Li et al., 2022a). Building on this success, the diffusion Transformer (Peebles & Xie, 2023; Fei et al., 2024d) and its variants (Bao et al., 2023; Fei et al., 2024b) have replaced the convolutional-based U-Net backbone (Ronneberger et al., 2015) with Transformers, resulting in greater scalability and more straightforward parameter expansion compared to U-Net diffusion models. This scalability advantage has been particularly evident in domains such as video generation (Ma et al., 2024b), image generation (Chen et al., 2023), and speech generation (Liu et al., 2023). Notably, recent works such as Make-an-audio 2 (Huang et al., 2023c, a) and StableAudio 2 (Evans et al., 2024) also explored the DiT architecture for audio and sound generation. In contrast, our work investigates the effectiveness of new multi-modal diffusion Transformer structure similar to Flux and optimized it with rectified flow.
@@ 3.1 Rectified Flow Trajectories
where the conditional vector fields $u_{t}(z|\epsilon)$ provides a tractable and equivalent objective. Although there exists different variants of the above formalism, we focus on Rectified Flows (RF) (Liu et al., 2022a; Albergo & Vanden-Eijnden, 2022; Lipman et al., 2023), which define the forward process as straight paths between the data distribution and a standard normal distribution:
@@ 4.1 Experimental settings
We utilize the last hidden state of FLAN-T5-XXL as fine-grained textual information and the pooler output of CLAP-L as coarse textual features. Referring to (Liu et al., 2024), our training process involves 10-second music clips, randomly sampled from full tracks. The training configuration includes a batch size of 128, a gradient clipping threshold of 1.0, and a learning rate of 1e-4. During inference, we apply a rectified flow with 50 steps and use a guidance scale of 3.5. To ensure model stability and performance, we maintain a secondary copy of the model weights, updated every 100 training batches through an exponential moving average (EMA) with a decay rate of 0.99, following the approach outlined by Peebles & Xie (2023). For unconditional diffusion guidance, we independently set the outputs of each of the two text encoders to null with a probability of 10%.

### v2

@@ 1 Introduction
Music, as a form of artistic expression, holds profound cultural importance and resonates deeply with human experiences (Briot et al., 2017). The task of text-to-music generation, which involves converting textual descriptions of emotions, styles, instruments, and other musical elements into audio, offers innovative tools and new avenues for multimedia creation (Huang et al., 2023b). Recent advancements in generative models have led to significant progress in this area (Yang et al., 2017; Dong et al., 2018; Mittal et al., 2021). Traditionally, approaches to text-to-music generation have relied on either language models or diffusion models to represent quantized waveforms or spectral features (Agostinelli et al., 2023; Lam et al., 2024; Liu et al., 2024; Evans et al., 2024; Schneider et al., 2024; Fei et al., 2024a, 2023c; Chen et al., 2024b). Among these, diffusion models (Song et al., 2020), trained to reverse the process of data transformation from structured states to random noise (Sohl-Dickstein et al., 2015; Song and Ermon, 2020), have shown exceptional effectiveness in modeling high-dimensional perceptual data, including music (Ho et al., 2020; Huang et al., 2023c; Ho et al., 2022; Kong et al., 2020b; Rombach et al., 2022).
@@
Given the iterative nature of diffusion process, coupled with the significant computational costs and extended sampling times during inference, there has been a growing body of research focused on developing more efficient training strategy and accelerating sampling schedule  (Karras et al., 2023; Liu et al., 2022a; Lu et al., 2022a; Fei, 2019; Lu et al., 2022b; Kingma and Gao, 2024), such as distillation (Sauer et al., 2024; Song et al., 2023; Song and Dhariwal, 2023). A particularly effective approach involves defining a forward path from data to noise, which facilitates more efficient training (Ma et al., 2024a) as well as better generative performance. One effective method among them is the Rectified Flow (Liu et al., 2022a; Albergo and Vanden-Eijnden, 2022; Lipman et al., 2023), where data and noise are connected along a linear trajectory. It offers improved theoretical properties and has shown promising results in image generation(Ma et al., 2024a; Esser et al., 2024), however, its application in music creation remains largely unexplored.
@@
In the design of model architectures, traditional diffusion models frequently employ U-Net (Ronneberger et al., 2015) as the foundational structure. However, the inherent inductive biases of convolutional neural networks inadequately captures the spatial correlations within signals (Esser et al., 2021) and are insensitive to scaling laws (Li et al., 2024b). Transformer-based diffusion models have effectively addressed these limitations (Peebles and Xie, 2023; Bao et al., 2023; Fei et al., 2024b, c) by treating images as sequences of concatenated patches and utilizing stacked transformer blocks for noise prediction. The incorporation of cross-attention for integrating textual information has established this approach as the standard for generating high-resolution images and videos from natural language descriptions, demonstrating impressive generalization capabilities (Chen et al., 2023, 2024a; Fei et al., 2024e; Esser et al., 2024; Fei et al., 2023b; Ma et al., 2024b, b; Yang et al., 2024). Notably, the recently open-sourced FLUX model, with its well-designed structure, exhibits strong semantic understanding and produces high-quality images, positioning it as a promising framework for conditional generation tasks.
@@ 2.1 Text-to-music Generation
Text-to-music generation seeks to produce music clips that correspond to descriptive or summarized text inputs. Prior approaches have primarily employed language models (LMs) or diffusion models (DMs) to generate quantized waveform representations or spectral features. For generating discrete representation of waveform, models such as MusicLM (Agostinelli et al., 2023), MusicGen (Copet et al., 2024), MeLoDy (Lam et al., 2024), and JEN-1 (Li et al., 2024c) utilize LMs and DMs on residual codebooks derived from quantization-based audio codecs (Zeghidour et al., 2021; Défossez et al., 2022). Conversely, models like Moûsai (Schneider et al., 2024), Noise2Music (Huang et al., 2023b), Riffusion (Forsgren and Martiros, 2022), AudioLDM 2 (Liu et al., 2024), MusicLDM (Chen et al., 2024b), and StableAudio (Evans et al., 2024) employ U-Net-based diffusion techniques to model mel-spectrograms or latent representations obtained through pretrained VAEs, subsequently converting them into audio waveforms using pretrained vocoders (Kong et al., 2020a). Additionally, models such as Mustango (Melechovsky et al., 2023) and Music Controlnet (Wu et al., 2024) incorporate control signals or personalization (Plitsis et al., 2024; Fei et al., 2023a), including chords and beats, in a manner similar to ControlNet (Zhang et al., 2023). Our method along with this approach by modeling the mel-spectrogram within a latent VAE space.
@@ 2.2 Diffusion Transformers
Transformer architecture (Vaswani et al., 2017) has achieved remarkable success in language models (Radford, 2018; Radford et al., 2019; Raffel et al., 2020) and has also demonstrated significant potential across various computer vision tasks, including image classification (DOSOVITSKIY, 2020; He et al., 2022; Touvron et al., 2021; Zhou et al., 2021; Yuan et al., 2021; Han et al., 2021), object detection (Liu et al., 2021; Wang et al., 2021, 2022; Carion et al., 2020), and semantic segmentation (Zheng et al., 2021; Xie et al., 2021; Strudel et al., 2021), among others (Sun et al., 2020; Li et al., 2022b; Zhao et al., 2021; Liu et al., 2022b; He et al., 2022; Li et al., 2022a). Building on this success, the diffusion Transformer (Peebles and Xie, 2023; Fei et al., 2024d) and its variants (Bao et al., 2023; Fei et al., 2024b) have replaced the convolutional-based U-Net backbone (Ronneberger et al., 2015) with Transformers, resulting in greater scalability and more straightforward parameter expansion compared to U-Net diffusion models. This scalability advantage has been particularly evident in domains such as video generation (Ma et al., 2024b), image generation (Chen et al., 2023), and speech generation (Liu et al., 2023). Notably, recent works such as Make-an-audio 2 (Huang et al., 2023c, a) and StableAudio 2 (Evans et al., 2024) also explored the DiT architecture for audio and sound generation. In contrast, our work investigates the effectiveness of new multi-modal diffusion Transformer structure similar to Flux and optimized it with rectified flow.
@@ 3.1 Rectified Flow Trajectories
where the conditional vector fields $u_{t}(z|\epsilon)$ provides a tractable and equivalent objective. Although there exists different variants of the above formalism, we focus on Rectified Flows (RF) (Liu et al., 2022a; Albergo and Vanden-Eijnden, 2022; Lipman et al., 2023), which define the forward process as straight paths between the data distribution and a standard normal distribution:
@@ 4.1 Experimental settings
We utilize the last hidden state of FLAN-T5-XXL as fine-grained textual information and the pooler output of CLAP-L as coarse textual features. Referring to (Liu et al., 2024), our training process involves 10-second music clips, randomly sampled from full tracks. The training configuration includes a batch size of 128, a gradient clipping threshold of 1.0, and a learning rate of 1e-4. During inference, we apply a rectified flow with 50 steps and use a guidance scale of 3.5. To ensure model stability and performance, we maintain a secondary copy of the model weights, updated every 100 training batches through an exponential moving average (EMA) with a decay rate of 0.99, following the approach outlined by Peebles and Xie (2023). For unconditional diffusion guidance, we independently set the outputs of each of the two text encoders to null with a probability of 10%.